# Event-Based Tracking Any Point with Motion-Augmented Temporal Consistency

### v1

@@ 4 Experiment
Figure 4: Some examples from the Ev-PointOdyssey dataset. Each example displays the RGB image in the top-left corner, with the event modality visualization in the bottom-right corner.
Table 1: The performance of the evaluated trackers on the Ev-PointOdyssey dataset are reported in terms of $\sigma_{avg}$, MTE, $Survival_{50}$. $\sigma_{avg}$ reflects the proportion of the error between the predicted and the ground truth within a certain range. MTE represents the error between the predicted trajectory and the ground truth. $Survival_{50}$ indicates the duration of tracking. ”Dark red” and ”Orange” represent feature tracking and optical flow models. Best results are in bold; second-best are underlined.
  Methods Modality Ev-PointOdyssey Params FPS
$\sigma_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  PIPs [15] Video 0.273 0.640 0.423 28.7M 119.1
PIPs++ [38] 0.336 0.270 0.505 17.6M 122.5
TAPIR [10] 0.322 0.515 0.446 29.3M 153.2
Context-PIPs [36] 0.331 0.630 0.491 30.5M 146.1
EKLT [11] Event 0.254 0.842 0.174 \ \
DeepEvT [25] 0.263 0.764 0.231 185.9M 158.6
E-RAFT [12] Event 0.265 0.789 0.176 5.3M 125.4
B-FLOW [13] 0.271 0.683 0.195 5.9M 78.4
Ours Event 0.358 0.262 0.553 6.6M 239.1
Table 2: The performance of different point tracking methods on the EC and EDS datasets. The proposed approach achieves the best performance on both datasets, with particularly notable improvements on the EDS dataset, which involves more camera motion.
  Methods EDS EC
@@
Ours 0.616 0.529 0.854 0.834
Dataset. To validate the effectiveness of the proposed method, this paper simulates events from the PointOdyssey dataset [38], referred to here as Ev-PointOdyssey. Compared to previous datasets [9, 15], PointOdyssey offers longer durations and more annotated points on average. In practice, event generation requires a continuous visual signal, which is typically achieved by rendering high-frame-rate videos for seamless flow. Here, the method from [29] is applied to minimize pixel displacement between frames. Subsequently, DVS-voltemeter [23] is utilized to synthesize realistic event data. Some examples from the Ev-PointOdyssey dataset are shown in Fig. 4.
To evaluate performance on real-world data, experiments are also conducted on the Event Camera (EC) dataset [27] and Event-aided Direct Sparse Odometry (EDS) dataset [17]. The EC dataset includes 240×180 resolution event streams and videos recorded with a DAVIS240C sensor. The EDS dataset contains videos and events captured simultaneously using a beam splitter, with the event data recorded at 640×480 resolution by the Prophesee Gen 3.1 sensor.
Metrics. For Ev-PointOdyssey, this paper follows the experimental setup of [38], using $\sigma_{avg}$, MTE, and $Survival_{50}$ for evaluation. $\sigma_{avg}$ measures the percentage of trajectories within error thresholds of $\{1,2,4,8,16\}$, averaged across these values. Median Trajectory Error (MTE) calculates the distance between predicted and the ground truth, using the median to reduce the impact of outliers. $Survival_{50}$ indicates the average duration until tracking failure, expressed as a percentage of the total sequence length. Failure is defined as an L2 distance exceeding $50$ pixels. For fair comparison, the temporal window for TS is aligned with the ground-truth time resolution. Additionally, model parameters and inference speed are reported for comparing the resource demands of different methods.
The EC and EDS datasets, which are commonly used to evaluate event-based feature tracking methods, provide ground truth for feature points. Quantitative metrics follow the setup of [25], utilizing Feature Age (FA) and Expected Feature Age (EFA). FA measures the percentage of successful tracking steps across thresholds from $1$ to $31$, with the final score being the average across all thresholds. EFA quantifies the lost tracks by calculating the ratio of stable tracks to ground truth and scaling it by the feature age.
@@
Implementation details. The model is trained on the Ev-PointOdyssey dataset with event clips at a spatial resolution of $256\times 320$ and a temporal length of $1.6$ seconds, optimized using Mean Absolute Error (MAE) loss. The AdamW optimizer [24] and OneCycleLR scheduler [32] are applied with a maximum learning rate of $45e-$ 4 and a cycle percentage of $0.1$. Subsequently, the model is fine-tuned with the same temporal length but a higher resolution of $512\times 640$. Experiments are conducted in parallel on 4 Nvidia RTX A6000 GPUs, implemented in PyTorch.
4.1 Quantitative Comparison
Baselines. The proposed method is compared with video-based approaches for TAP such as PIPs [15], PIPs++ [38], TAPIR [10], and Context-PIPs [36] to validate the advantages of the event modality in this task. In the absence of event-based approaches for TAP, we extend several state-of-the-art (SOTA) point correspondence methods based on events. EKLT [11] and DeepEvT [25] are event-based feature tracking methods. EKLT is built on first principles of events and can be directly applied to this dataset, while DeepEvT is retrained on the Ev-PointOdyssey dataset. E-RAFT [12] and B-FLOW [13] serve as event-based optical flow estimation methods. We link the optical flow across times, using bilinear interpolation to calculate flow for sub-pixel points. Predicted coordinates for points that exceed the boundaries are clamped to the boundary.
(a) Translation
(b) Rotation
Figure 5: The two figures present the results of dense point tracking by the proposed method on the EC dataset. (a) Camera translation sequence. (b) Camera rotation sequence.
(a) PIPs++ [38]
(b) E-RAFT [12]
(c) DeepEvT [25]
(d) Ours
(e) Ground Truth
Figure 6: Qualitative results for EDS (top two rows), Ev-PointOdyssey (bottom two rows). The predicted trajectories are visualized using a pink-to-yellow colormap, with sparse ground truth marked by “$\times$”. The first row highlights the tracking performance of different methods in low-texture regions (white tabletop), where DeepEvT almost loses tracking. In the second row, the target point (table corner) moves out of view and then returns, causing E-RAFT to fail because it only models inter-frame pixel displacement. The last two rows show dense tracking results on the Ev-PointOdyssey dataset, highlighting the superiority of the proposed method.
Ev-PointOdyssey results. On the Ev-PointOdyssey dataset, the proposed method outperforms video-based methods across all three metrics, as shown in Tab. 1. Video-based methods rely on iterative local searches to match appearance information, which leads to errors when faced with large displacements or nonlinear motion. In contrast, the proposed method incorporates kinematic features to guide matching, effectively addressing these challenges. Compared to event-based methods, EKLT and DeepEvT are designed for feature tracking as they directly predict the displacement of tracking points from local event streams. However, they struggle in low-texture areas due to insufficient event data, causing tracking failures. The proposed method integrates spatial context through local-global iterative matching, allowing tracking in low-texture regions. E-RAFT and B-FLOW estimate pixel displacement over short intervals but are susceptible to disruption from occlusions or points moving out of bounds. Although this paper does not explicitly model occlusions, its ability to capture long-term motion dependencies enables it to maintain stable tracking even when points become temporarily invisible. Moreover, the proposed method is parameter-efficient and provides faster runtime. This efficiency stems from two factors: first, event cameras focus solely on dynamic changes, thereby reducing visual redundancy; second, a transformer replaces the MLPs [38, 36] to effectively capture temporal dependencies. As EKLT is not a deep learning algorithm, a fair comparison of its parameter count and FPS is not feasible, therefore it is not included in the table.
EC and EDS results. Similar to the results on Ev-PointOdyssey, the proposed method outperforms existing event-based trackers on real-world datasets, as reported in Tab. 2. The EDS dataset, with faster camera motion than the EC dataset, results in generally lower performance across all methods. Nevertheless, the proposed method effectively handles the noise introduced by this, ensuring stable tracking results. While these metrics reflect feature tracking performance, the proposed approach also demonstrates superior performance on TAP. Figure 5 presents two sequences from the EC dataset: one with translational camera motion and the other with rotational camera motion, highlighting the robustness of the proposed method in tracking dense points across diverse motion patterns.
4.2 Qualitative Analysis
Figure 6 illustrates a comparison of the proposed method with prior works on the EDS and Ev-PointOdyssey datasets. PIPs++ takes a sequence of 48 RGB frames as input, while other methods rely on event data corresponding to the same time period. Trajectories are shown in a pink-to-yellow colormap, indicating point movement from the pink starting position to the yellow endpoint. Due to space constraints, 48 ground truth points are downsampled to 9, marked with an “$\times$” pattern in the first two rows.
The top two rows show the motion of the tabletop (a low-texture point) and the table corner (a key point). The tabletop point remains consistently visible, while the table corner temporarily exits the field of view before reentering. The first row depicts the predicted trajectory of the tabletop point across different methods. Here, DeepEvT struggles due to its reliance on local event data, which limits accuracy in low-texture areas with sparse events, leading to missed points. The second row highlights the predicted trajectory of the table corner point. E-RAFT fails to maintain tracking when points move out of the frame because it models only short-term pixel displacements rather than the motion of physical surface points. For both types of points, PIPs++ exhibits lower tracking accuracy under rapid motion compared to our method, particularly on EDS, which involves intense camera movement. The bottom two rows display dense tracking results on Ev-PointOdyssey, further demonstrating the superiority of the proposed approach.
@@ 4.2 Qualitative Analysis
       MGM VMAM Ev-PointOdyssey
PF MLP CA TC TA $\sigma_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  0.324 0.385 0.485
0.335 0.367 0.489
0.340 0.326 0.495
0.349 0.291 0.510
0.348 0.288 0.506
0.358 0.262 0.553
(a) Kinematic vectors w/o MLP
(b) Kinematic vectors w/ MLP
Figure 7: Impact of the MLP in MGM. Arrows represent the direction of kinematic vectors, and their length indicates intensity.
@@ 4.3 Ablation Study
Impact of the motion-guidance module. The first three rows of Tab. 3 illustrate how kinematic features from MGM contribute to point displacement updates, where PF stands for Plane Fitting. The first row shows a baseline model without MGM, relying solely on appearance feature matching. The second row includes MGM but omits MLP-based correction for ambiguous kinematic features. Results indicate that kinematic features play a significant role in enhancing tracking accuracy, and this effect is further amplified when corrected by the MLP. As shown in Fig. 7, the initial kinematic features are inaccurate at some time steps due to object motion overlap, but they exhibit temporal coherence after MLP correction.
Effectiveness of variable motion aware module. Rows four through six in Tab. 3 present the ablation studies for each component of VMAM, where CA, TC, and TA represent Cross Attention, Temporal Convolution, and Temporal Attention, respectively. The fourth row performs correlations solely within appearance features, without guidance from kinematic features. The fifth row employs 1D temporal convolutions to capture short-term dependencies. In the final row, the initial experimental setup is maintained, modeling temporal relationships via both short-term and long-term paths. The results reveal that incorporating both kinematic guidance and temporal modeling progressively enhances performance, highlighting the effectiveness of each VMAM component. Figure 8 provides a visual comparison. Features without VMAM exhibit temporal discontinuities, while VMAM-enhanced features demonstrate improved temporal consistency with lower standard deviation.
@@
Figure 8: Effectiveness of VMAM. The first row shows results using only appearance features, without VMAM, while the second row includes VMAM. (a) Temporal distribution across feature channels. (b) Temporal standard deviation across feature channels.
@@ 5 Conclusion
This study introduces a novel event-based framework for tracking any point, leveraging a motion-guidance module to extract kinematic features that refine the matching process and constructs a dynamic-appearance space. Additionally, the integration of a variable motion aware module enables the system to account for motion variations, ensuring temporal consistency across diverse velocities. To evaluate the approach, a simulated event point tracking dataset was collected. The proposed method achieved state-of-the-art tracking performance on both the simulated dataset and two real-world datasets, with competitive model parameters and faster inference time. This technique holds substantial potential for applications in embodied intelligence, autonomous driving, and related fields.

### v2

@@ 4 Experiment
Figure 4: Some examples from the Ev-PointOdyssey (top row) and Ev-Davis (bottom row) dataset. Each example displays the RGB image in the top-left corner, with the event modality visualization in the bottom-right corner.
Table 1: The performance of the evaluated trackers on the Ev-PointOdyssey and Ev-davis datasets are reported in terms of $\delta_{avg}$, MTE and $Survival$. $\delta_{avg}$ reflects the proportion of the error between the predicted and the ground truth within a certain range. MTE represents the error between the predicted trajectory and the ground truth. $Survival$ indicates the duration of tracking. “Ev-” indicates that video input is replaced with event data, and the model is retrained on Ev-PointOdyssey. Best results are in bold; second-best are underlined.
  Method Modality Ev-PointOdyssey Ev-Davis Params
$\delta_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$ $\delta_{avg}\uparrow$ MTE $\downarrow$ $Survival_{16}\uparrow$
  PIPs++ [44] Video 0.336 26.95 0.505 0.697 4.16 0.897 17.6M
Cotracker [21] 0.302 - 0.552 0.726 3.99 0.902 24.2M
Cotracker3 [20] 0.364 26.84 0.409 0.739 3.73 0.913 25.4M
Ev-PIPs++ [44] Event 0.306 57.36 0.475 0.668 5.21 0.874 24.2M
Ev-Cotracker3 [21] 0.316 42.55 0.428 0.696 4.52 0.887 25.4M
Ours Event 0.362 25.63 0.560 0.747 3.47 0.928 6.6M
Table 2: The performance of different feature tracking methods on the EC and EDS datasets. The proposed approach achieves the best performance on both datasets, with particularly notable improvements on the EDS dataset, which involves more camera motion. “E+V” represents events and videos.
  Method Input EDS EC
@@
FE-TAP [25] 0.676 0.589 0.844 0.838
EM-ICP [45] Event 0.161 0.120 0.337 0.334
HASTE [3] 0.096 0.063 0.442 0.427
AEB-Tracker [38] 0.473 0.466 0.553 0.473
ETAP [14] 0.701 0.601 0.881 0.876
Ours Event 0.713 0.626 0.885 0.875
Sec. 4.1 details the experimental setup. The next two sections comprehensively evaluate the proposed method on two tasks: tracking of any point and feature tracking. The key difference is that feature tracking focuses on high-contrast points using local descriptors, while tracking of any point extends this ability to low-texture regions. Due to the lack of prior work on event-based tracking of any point, comparisons are made with state-of-the-art video-based methods and their direct adaptation to the event domain. Feature tracking is compared with existing event-based methods and video-based TAP method adapted to this task, as feature tracking can essentially be seen as a specialized subset of the more general tracking any point problem.
4.1 Implementation Details
Event time surface is generated by normalizing event timestamps and encoding them into two separate channels for positive and negative polarity events. The temporal window length is set according to the ground truth frequency, and timestamps within each window remain untruncated to fully preserve the motion trajectory over time.
@@ 4.1 Implementation Details
The model is trained on the Ev-PointOdyssey dataset at a spatial resolution of 512×640512$\times$ 640512 × 640. Optimization is performed using the AdamW optimizer [26] and OneCycleLR scheduler [34] with a maximum learning rate of $5e-4$ and a cycle percentage of 0.10.10.1. Weighted Mean Absolute Error loss is used for supervision, defined as:
$L=\sum_{i}\gamma^{T-i}\cdot\lvert y_{pred}^{i}-y_{true}^{i}\lvert,i=1,\dots,T,$ (8)
where $\gamma$ is the weight factor, set to 0.8, and $T$ represents the number of bins per input sequence, set to 48. The purpose of this weighting is to impose a greater penalty on errors in later time steps, thereby improving the performance on long-term tracking. Experiments are conducted in parallel on 8 Nvidia RTX A6000 GPUs, implemented in PyTorch.
Figure 5: Qualitative results for Ev-PointOdyssey and Ev-Davis datasets. (a) The predicted trajectories are visualized using a blue-to-green colormap. By leveraging kinematic vectors for matching instead of relying solely on appearance features, the proposed method exhibits strong robustness to occlusion. (b) Due to missing ground truth for occluded points in the Ev-Davis dataset, only their positions are visualized. Notably, the proposed method leverages the high temporal resolution of event cameras to maintain stable tracking during fast motion (see red and yellow points), while video-based methods suffer from motion blur, leading to performance degradation.
4.2 Task1: TAP
Dataset. To validate the effectiveness of the proposed method, this paper simulates events from the PointOdyssey dataset [44] and TAP-Vid-Davis dataset [9], referred to as Ev-PointOdyssey and Ev-Davis. Compared to previous datasets [9, 16], PointOdyssey offers longer durations and more annotated points on average, with each video lasting approximately 2000 frames and 68 seconds on average, whereas prior datasets contain around 60 frames per video and last about 2 seconds. TAP-Vid-Davis is a real-world captured dataset, featuring faster motion and frequent point occlusions in the videos.
[…]
Quantitative results. Table 1 compares the proposed method with video-based approaches for TAP, such as PIPs++ [44], Cotracker [21] and Cotracker3 [20]. Additionally, due to the absence of publicly available code for event-based methods, some video-based methods are directly applied to the event domain for a fairer comparison. However, these video-based methods rely solely on local iterative appearance matching, making them prone to errors in large displacements or nonlinear motion. This issue is further amplified when applied to event data, as events lack texture information. In contrast, the proposed method integrates kinematic vectors to guide matching, improving robustness in these challenging scenarios. $Survival_{50}$ metric shows a 17.9% improvement over methods that solely rely on event data. In addition, it exhibits highly competitive performance even when compared to approaches that utilize video input. Moreover, the proposed method is parameter-efficient. This efficiency stems from two factors: first, event cameras focus solely on dynamic changes, thereby reducing visual redundancy and enabling content learning without complex networks; second, a transformer replaces MLPs [44, 39] to effectively capture temporal dependencies.
Qualitative analysis. Figure 5 illustrates a comparison of the proposed method with prior works on the Ev-PointOdyssey and Ev-Davis datasets. PIPs++ [44] and Cotracker3 [20] take video input, while the proposed method takes event input. Trajectories are shown in a blue-to-green colormap, indicating point movement. Due to the lack of ground truth for occluded points in the Ev-Davis dataset, accurately visualizing point trajectories is difficult, so only the point positions are visualized.
The left panel (a) shows results on Ev-PointOdyssey, where the query point is initially visible, gets occluded as motion starts, then reappears. Video-based methods like PIPs++ and Cotracker3, relying solely on appearance matching, struggle under these conditions. In contrast, the proposed method mitigates matching ambiguities caused by occlusion by using kinematic vectors as guidance. The right panel (b) shows results on the Ev-Davis, where motion blur degrades video-based methods. Leveraging the high temporal resolution of event cameras, the proposed method demonstrates robust tracking in this challenging scenario.
4.3 Task2: Feature Tracking
Dataset. The Event Camera (EC) dataset [30] and Event-aided Direct Sparse Odometry (EDS) dataset [18] are commonly used to evaluate event-based feature tracking methods.The EC dataset includes 240×180240$\times$ 180240 × 180 resolution event streams and videos recorded with a DAVIS240C sensor. The EDS dataset contains videos and events captured simultaneously using a beam splitter with spatial resolution 640×480640$\times$ 480640 × 480. Notably, both datasets provide asynchronous ground truth for feature points at varying timestamps. To ensure fair comparisons, spline interpolation is applied to align the ground truth to a common timestamp.
(a) DeepEvT [28]
(b) PIPs++ [44]
(c) Ours
(d) GT
Figure 6: Tracking results on EC (top row) and EDS (bottom row) dataset. DeepEvT fails to retain points under occlusion, while PIPs++ experiences performance degradation during rapid motion.
Metric. Quantitative metrics follow the setup of [27], utilizing Feature Age (FA) and Expected Feature Age (EFA). FA measures the percentage of successful tracking steps across thresholds from 111 to 313131, with the final score being the average across all thresholds. EFA quantifies the lost tracks by calculating the ratio of stable tracks to ground truth and scaling it by the feature age.
Results. The proposed method is compared against three categories of approaches across two datasets. First, video-based methods for tracking of any point, such as PIPs++ [44] and Cotracker3 [20]. On the EDS dataset, motion blur caused by rapid movement degrades their performance, while on the EC dataset, low resolution leads to texture blur, further affecting their accuracy. Second, methods that utilize both video frames and event data, including EKLT [11], DeepEvT [28], and FE-TAP [25]. Among these, FE-TAP integrates events and video for tracking of any point. Since its implementation is not publicly available, the results reported in the original paper are used for a fair comparison. Lastly, event-based tracking methods, such as EM-ICP [45], HASTE [3], AEB-Tracker [38] and ETAP [14], are also evaluated. Across both datasets, the proposed method outperforms all three categories of existing trackers, see Tab. 2. Figure 6 provides a visual comparison, showing that our method better maintains consistency under challenging conditions by leveraging motion cues and temporal modeling.
@@ 4.3 Task2: Feature Tracking
       MGM VMA Ev-PointOdyssey
PF MLP CA TC TA $\delta_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  0.312 43.64 0.488
0.325 38.72 0.485
0.340 35.65 0.513
0.360 29.56 0.538
0.358 27.63 0.549
0.362 25.63 0.560
(a) Query frame
(b) w/o MGM
(c) w/ MGM
Figure 7: Effectiveness of MGM. (a) The initial frame containing the query point. (b) Results from appearance-based matching without MGM. (c) Results incorporating kinematic vectors extracted by MGM for enhanced matching.
@@ 4.4 Ablation Study
Impact of the motion-guidance module. The first three rows of Tab. 3 illustrate how kinematic vectors from MGM contribute to point displacement updates, where PF stands for Plane Fitting. The first row shows a baseline without MGM, relying solely on appearance matching. The second row includes MGM but omits MLP-based correction for ambiguous kinematic vectors. Results indicate that kinematic vectors play a significant role in enhancing tracking accuracy, and this effect is further amplified when corrected by the MLP. As shown in Fig. 7, MGM is essential for handling low-texture regions, where appearance-based matching alone often fails due to the lack of texture information in event data. By introducing kinematic vectors as additional motion cues, MGM effectively compensates for this limitation, improving robustness in challenging scenarios.
Effectiveness of variable motion aware module. Rows four through six in Tab. 3 present the ablation studies for each component of VMA, where CA, TC, and TA represent Cross Attention, Temporal Convolution, and Temporal Attention, respectively. The fourth row performs appearance matching guided by kinematic vectors. The fifth row employs 1D temporal convolutions to capture short-term dependencies. The final row retains the full setup, modeling temporal relationships via both short- and long-term paths. Results show that integrating kinematic guidance and temporal modeling progressively enhances performance, demonstrating the effectiveness of each VMA component. Figure 8 offers a visual comparison. Features without CA, TC, and TA (i.e., $\mathbf{G}^{k}$) show temporal discontinuities, while enhanced features exhibit improved consistency and lower standard deviation.
@@
Figure 8: Effectiveness of VMA. The first row shows results using only appearance features, without VMA, while the second row includes VMA. (a) Temporal distribution across feature channels. (b) Temporal standard deviation across feature channels.
@@ 5 Conclusion
This study introduces a novel event-based framework for tracking any point, leveraging a motion-guidance module to extract kinematic features that refine the matching process and constructs a dynamic-appearance space. Additionally, the integration of a variable motion aware module enables the system to account for motion variations, ensuring temporal consistency across diverse velocities. To evaluate the approach, two simulated event point tracking dataset was collected. The proposed method outperforms existing approaches on both tracking of any point and feature point tracking tasks across four datasets, with competitive model parameters. This technique holds substantial potential for applications in embodied intelligence and related fields.
Acknowledgment
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008 and 62306295.
[…]
Supplementary Material
6 Network Structure
Network Encoder. The encoder is based on ResNet [17] and consists of four feature extraction modules, each containing two stacked residual blocks. The first block in each module performs downsampling, while the second maintains the resolution. Features at different resolutions from each module are aligned to a consistent resolution via interpolation and concatenated along the channel to form multi-scale representations. The concatenated features are then compressed along the channel using a $\times 1$ 1 convolution.
7 Implementation Details
[…]
Data augmentations. To improve model robustness, several data augmentation techniques are applied. Random pixel erasure is performed at all time steps, with varying erased regions to enhance the network’s ability to handle occlusions. The erased values are replaced by the mean of the original region. Additionally, random scale-up is employed to improve tracking across different object scales, with scale parameters for the next time step initialized from the previous step and adjusted with slight random perturbations. Spatial flipping and temporal reversal are also applied to further diversify the dataset.
8 Additional Ablation Experiments
In addition to the ablation studies reported in the main paper, experiments are conducted on the input event representation and the number of pixels used for plane fitting in the motion-guidance module. Due to time constraints, these experiments are performed on a baseline model, which excludes the variable motion aware module, and are conducted at a lower resolution of 192×256192$\times$ 256192 × 256.
Table 4: Ablation studies on different input event representations.
  Representations Ev-PointOdyssey
$\sigma_{avg}\uparrow$ MTE $\downarrow$ $Survival_{50}\uparrow$
  Event image 0.298 54.43 0.443
Voxel grid 0.312 51.24 0.457
Time surface 0.323 48.41 0.477
(a) Ablation study on $\gamma$
(b) Ablation study on iterations
Figure 9: The survival metric is maximized when gamma and iteration are set to 0.8 and 6, respectively.
Input representation. Different input event representations significantly impact event-based networks. To evaluate their effects on this task, various representations are used to train the baseline model, with the results presented in Tab. 4. Event image aggregates the number of events at each pixel over a fixed time window. Voxel grid [47] represents a spatiotemporal histogram, where each voxel corresponds to a specific pixel and time interval. Time surface [29] is a 2D map where each pixel stores the timestamp of the most recent event. As highlighted in the introduction of main paper, the TAP task relies on continuous motion information. Both event image and voxel grid quantize event timestamps, discarding dense temporal details, which limits their ability to capture smooth motion dynamics. In contrast, time surface retains rich temporal information, with each pixel encoding the motion history at its location, making it more suitable for encoding temporal motion cues effectively. As a result, time surface demonstrates superior performance compared with other representations.
Figure 10: Qualitative results for Ev-PointOdyssey dataset.
Figure 11: Qualitative results for EDS dataset.
Weight factor $\gamma$ in loss function. As described in Eq. (8) of the main paper, the losses across different time steps are weighted by a factor $\gamma$, assigning a higher penalty to losses occurring at later time steps. An ablation study is performed to identify the optimal $\gamma$. Figure 9(a) demonstrates that the $Survival_{50}$ is maximized when $\gamma$ equals 0.8.
The number of iterations. In addition to $\gamma$, the number of iterations also significantly affects tracking accuracy. While increasing the iteration count typically enhances accuracy, it also requires more computational resources. To find an optimal trade-off, an ablation study on the iteration count is performed. As illustrated in Fig. 9(b), when the iteration count reaches 6, further increases do not yield significant improvements in the metrics.
9 More Qualitative Results
Additional qualitative results are provided in Figs. 10 and 11 with higher-resolution images for clearer visualization.